Mpho’s monthly disposable income increases from R1 800 to R2 300. As a result, his monthly savings increase from R290 to R440. This implies that his marginal propensity to consume is the following 


1
Expert's answer
2020-10-19T13:37:54-0400

"\bold {Answer}"

"MPC = \bold {0.7}"


"\bold {Solution}"

Marginal propensity to consume (MPC) measures the change in consumption derived from a dollar change in disposable income.

"MPC = \dfrac {∆C}{∆Y_D}" Where "\u2206C" is the change in consumption and "\u2206Y_D" is the change in disposable income.

Because income is either consumed or saved,

"C + S = Y_D" Where "C" represents consumption, "S" represents savings, and "Y_D" is the disposable income.


Consequently,

"MPC + MPS = 1," where "MPS" is the marginal propensity to save.

"=> MPC = 1-MPS"

MPS measures the change in savings derived from a dollar change in disposable income.

"MPS = \dfrac {∆S}{∆Y_D}"

"= \dfrac {R440-R290}{R2 \space 300 - R1 \space 800}"


"= \dfrac {R150}{R500}"


"= \bold {0.3}"


"=> MPC = 1-0.3"

"= \bold {0.7}"